# An Explicit Entire Function of Order One with All Zeros on a Line and Bounded in a Half-Plane

### v1

@@ Abstract.
We construct a single explicit entire function $\Xi_{c}(s)$ of order 1, with all zeros provably on $\operatorname{Re}(s)=1/2$, satisfying a functional equation $\Xi_{c}(s)=\Xi_{c}(1-s)$, whose normalized form $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$ is uniformly bounded for $\operatorname{Re}(s)>1+\delta$ (any $\delta>0$) yet satisfies $\sup_{t}|Z_{c}(1+it)|=+\infty$. The function thus satisfies an analogue of the Riemann Hypothesis together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$. The construction uses Hadamard products with controlled zero perturbations; the transition is proved unconditionally via a dyadic large-sieve argument and a Phragmén–Lindelöf transfer.
@@ 1. Introduction
  • On $\sigma=1$: $\limsup_{t\to\infty}|\zeta(1+it)|=\infty$, with $|\zeta(1+it)|=\Omega(\log\log t)$ [14, Thm. 8.5].
@@
This paper provides such a construction: an explicit entire function $\Xi_{c}(s)$ of order $1$ with all zeros provably on $\operatorname{Re}(s)=1/2$, satisfying $\Xi_{c}(s)=\Xi_{c}(1-s)$, whose normalized form is uniformly bounded for $\sigma>1$ yet unbounded on $\sigma=1$—the same sharp dichotomy exhibited by $\zeta$. The associated ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product, providing a concrete “RH-satisfying” $L$-function–like object whose analytic properties can be studied unconditionally.
@@
For any $c>0$, define $\Xi_{c}(s)$ as in §2, with $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$. Then:
@@
    For each $\delta>0$, $\sup_{t}|Z_{c}(1+\delta+it)|\leq\exp(c\,\zeta_{\alpha}(1+\delta)+O(1))$.
@@
Parts (a)–(c) hold by construction; (d) via the spectral identity (Proposition 3.1) with linearization error controlled by Lemma 1.2; and (e) via a Phragmén–Lindelöf argument (§5). Moreover, the ratio $Z_{c}/Z_{0}$ is itself a Dirichlet series with generalized Euler product, bounded for $\sigma>1$ and unbounded on $\sigma=1$, with all zeros provably on $\sigma=1/2$.
Lemma 1.2 (Linearization convergence).
(1) $\sum_{n=1}^{\infty}\frac{E(\gamma_{n}^{(0)})^{2}}{n^{2}}<\infty,$
where $E$ is the full (infinite) series (3) evaluated at the regular zero positions.
Proof.
Head (large sieve). $E_{K_{j}}$ is a trigonometric polynomial with $K_{j}$ terms and frequency range $\Lambda_{j}=\omega_{K_{j}}-\omega_{1}=O(j)$. The zeros $\gamma_{n}^{(0)}$ have spacing $\delta\geq c_{1}/j$ for $n\leq 2^{j+1}$. By the large sieve inequality for trigonometric polynomials at well-spaced points (cf. [10, Thm. 1], [11, §7]):
$\sum_{n\leq 2^{j+1}}\!|E_{K_{j}}(\gamma_{n}^{(0)})|^{2}\leq(\pi/\delta+\Lambda_{j})\,\tfrac{1}{\pi^{2}}\sum_{k\leq K_{j}}\alpha_{k}^{-1}\leq C_{1}\,j\log j,$
since $\sum_{k\leq K_{j}}\alpha_{k}^{-1}\sim\log\log p_{K_{j}}\sim\log j$ (Mertens). Therefore $\text{HEAD}_{j}\leq C_{1}\,j\log j\,/\,4^{j}$, and $\sum_{j}j\log j/4^{j}<\infty$.
Tail (Abel summation). From the proof of Lemma 2.1, the tail $D_{K}=\sum_{k>K}(\omega_{k+1}-\omega_{k})/\alpha_{k}^{1/2}$ is a left Riemann sum for the decreasing function $e^{-\omega/2}$. By Bertrand’s postulate (mesh width $\leq\log 2$): $D_{K}\leq\sqrt{2}\int_{\omega_{K}}^{\infty}e^{-\omega/2}\,d\omega=2\sqrt{2}\,\alpha_{K}^{-1/2}\leq 2\sqrt{2}/\sqrt{p_{K}}$. Hence $|R_{K_{j}}(\gamma_{n}^{(0)})|\leq(C_{\theta}/\pi)(\alpha_{K_{j}+1}^{-1/2}+\gamma_{n}D_{K_{j}})\leq C_{2}\,2^{j(1-\varepsilon)/2}/j^{3/2}$ (with $\varepsilon=0.1$, using $\gamma_{n}\leq C_{3}\,2^{j}/j$ and $\sqrt{p_{K_{j}}}\geq C_{4}\,2^{(1+\varepsilon)j/2}\sqrt{j}$). Therefore $\text{TAIL}_{j}\leq C_{5}/(j^{3}\cdot 2^{\varepsilon j})$, and $\sum_{j}1/(j^{3}\cdot 2^{\varepsilon j})<\infty$.
Both contributions decay exponentially, so $\sum S_{j}<\infty$. ∎
1.1. Method
Given a sequence of positive reals $0<t_{1},t_{2},\ldots$ with $\sum t_{n}^{-2}<\infty$, define the Hadamard product
(2) $\Xi(s)=\prod_{n=1}^{\infty}\left(1+\frac{(s-\tfrac{1}{2})^{2}}{t_{n}^{2}}\right).$
This has zeros at $s=1/2\pm it_{n}$ and satisfies $\Xi(s)=\Xi(1-s)$ (cf. [5, §1.10], [14, §2.12]). We choose the zero positions $\{t_{n}\}$ so that the counting function $N_{c}(T)=\#\{n:t_{n}\leq T\}$ approximates $N_{\mathrm{main}}(T)+cE(T)$, where $E$ is a trigonometric series whose Fourier amplitudes decay like the prime reciprocals.
The growth of $|Z|$ on vertical lines is controlled by the Poisson extension of $E$: bounded for width $w>1/2$ (i.e., $\sigma>1$) and unbounded at width $w=1/2$ (i.e., $\sigma=1$), from a marginal convergence/divergence condition on the damped Fourier amplitudes.
2. The Explicit Construction
2.1. Parameters
@@
Fix an irrational $\theta$ (e.g., $\theta=e^{\sqrt{2}}$), an amplitude $c>0$, and the primes $p_{1}<p_{2}<\cdots$. Define:
(3) $\alpha_{k}=p_{k},\quad\omega_{k}=\log p_{k},\quad\phi_{k}=k\theta,\quad E(T)=-\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(\omega_{k}T+\phi_{k})}{p_{k}^{1/2}}.$
Lemma 2.1 (Convergence).
The series $E(T)$ converges for all $T>0$.
Proof.
Write $E(T)=-(1/\pi)\operatorname{Im}\bigl[\sum_{k}d_{k}\,c_{k}\bigr]$ where $c_{k}=e^{ik\theta}$ and $d_{k}=e^{i\omega_{k}T}/\alpha_{k}^{1/2}$. Apply summation by parts (Abel’s formula): $\sum_{k=1}^{K}c_{k}d_{k}=G_{K}d_{K}-\sum_{k=1}^{K-1}G_{k}(d_{k+1}-d_{k})$, where $G_{k}=\sum_{j=1}^{k}e^{ij\theta}$ satisfies $|G_{k}|\leq 2/|1-e^{i\theta}|=:C_{\theta}$ (since $\theta$ is irrational). The boundary term satisfies $|G_{K}d_{K}|\leq C_{\theta}\,\alpha_{K}^{-1/2}\to 0$. For the variation sum: $|d_{k+1}-d_{k}|\leq|\omega_{k+1}-\omega_{k}|\,T/\alpha_{k}^{1/2}+(\alpha_{k}^{-1/2}-\alpha_{k+1}^{-1/2})$. The amplitude differences telescope: $\sum(\alpha_{k}^{-1/2}-\alpha_{k+1}^{-1/2})=\alpha_{1}^{-1/2}$. The phase differences satisfy $\sum_{k}(\omega_{k+1}-\omega_{k})/\alpha_{k}^{1/2}=\sum_{k}\log(\alpha_{k+1}/\alpha_{k})/\alpha_{k}^{1/2}<\infty$ (this sum is a Riemann sum for $\int_{\omega_{1}}^{\infty}e^{-\omega/2}\,d\omega<\infty$, with bounded mesh by Bertrand’s postulate). Hence $\sum|G_{k}(d_{k+1}-d_{k})|$ converges absolutely, and $G_{K}d_{K}\to 0$, so $\sum c_{k}d_{k}$ converges. ∎
@@ 2.2. The perturbed zeros
and let $\Xi_{c}(s)$ be the Hadamard product (2) with $t_{n}=\gamma_{n}$.
Lemma 2.2 (Well-definedness of the Hadamard product).
For any $c>0$, the sequence $\{\gamma_{n}\}$ satisfies $\gamma_{n}>0$ for all $n$ (taking $c$ small enough for finitely many small $n$) and $\sum\gamma_{n}^{-2}<\infty$, so $\Xi_{c}(s)$ is an entire function of order $1$.
@@
Since $E(T)$ converges for each $T$ (Lemma 2.1), $\delta_{n}$ is finite for each $n$. For $n$ large, $|\delta_{n}|=O(c|E(\gamma_{n}^{(0)})|/\log\gamma_{n}^{(0)})$. By the Abel-summation bound of Lemma 2.1, $|E(T)|=O(T)$ for each $T$ (the bound grows at most linearly since the total variation of $e^{i\omega_{k}T}$ over $k\leq K$ is $O(T\log\alpha_{K})$, controlled by the monotone decay of $\alpha_{k}^{-1/2}$). Thus $|\delta_{n}|=O(c\gamma_{n}^{(0)}/\log\gamma_{n}^{(0)})<\gamma_{n}^{(0)}/2$ for $n$ sufficiently large, giving $\gamma_{n}>\gamma_{n}^{(0)}/2>0$. Since $\gamma_{n}^{(0)}\sim 2\pi n/\log n$, we have $\sum\gamma_{n}^{-2}\leq 4\sum(\gamma_{n}^{(0)})^{-2}<\infty$. The exponent of convergence equals $1$. ∎
Remark 2.3.
The zeros $\gamma_{n}$ need not be in increasing order: crossings ($\gamma_{n+1}<\gamma_{n}$) may occur where $|E^{\prime}|$ is large. For a finite truncation $E_{K}$, global monotonicity of $F_{K}(T)=N_{\mathrm{main}}(T)+cE_{K}(T)$ holds for $c<\log(T_{0}/(2\pi))/(2\pi C_{K})$ with $C_{K}=\frac{1}{\pi}\sum_{k=1}^{K}\omega_{k}\alpha_{k}^{-1/2}$. For the infinite series, occasional crossings produce zeros of multiplicity >1>1 without affecting the Hadamard product convergence or the spectral identity (which depends only on the linearization at $\gamma_{n}^{(0)}$, not on the ordering of $\gamma_{n}$).
@@ 3. The Spectral Identity
The following proposition relates the Hadamard product ratio $\Xi_{c}/\Xi_{0}$ to the Fourier content of $E$. All quantities ($\alpha_{k},\omega_{k},\phi_{k},E$) are as defined in §2.
@@
For $\sigma\geq 1$ and $|t|\geq 2$:
(6) $\log\frac{|\Xi_{c}(\sigma+it)|}{|\Xi_{0}(\sigma+it)|}=-c\sum_{k}\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c).$
@@
Step 1 (Linearization). By (5), $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$. Each Hadamard factor satisfies (Taylor-expanding in $\delta_{n}$):
@@
Step 2 (Finite-truncation limit). For each $K$, let $\Xi_{c,K}$ be the Hadamard product with zeros defined using the truncation $E_{K}$ (first $K$ terms of $E$). The perturbations $\delta_{n,K}=-cE_{K}(\gamma_{n}^{(0)})/N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})$ satisfy $|\delta_{n,K}|\leq 2\pi cM_{K}/\log\gamma_{n}^{(0)}$ where $M_{K}=\frac{1}{\pi}\sum_{k=1}^{K}\alpha_{k}^{-1/2}$ is a constant. The linearization error for the $K$-truncation is then
(7) $\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\int_{\gamma_{1}^{(0)}}^{\infty}\frac{2w^{2}\,E_{K}(u)}{u(u^{2}+w^{2})}\,du+O(c^{2}),$
where $O(c^{2})$ accounts for the quadratic remainder $\sum\delta_{n,K}^{2}/(\gamma_{n}^{(0)})^{2}$, which converges since $|\delta_{n,K}|=O(cM_{K}/\log\gamma_{n}^{(0)})$ and $\sum 1/(\gamma_{n}^{(0)}\log\gamma_{n}^{(0)})^{2}<\infty$. This error is $O(c^{2}M_{K}^{2})$, independent of $t$.
Since $E_{K}\to E$ pointwise and $\gamma_{n,K}\to\gamma_{n}$ for each $n$, the Hadamard products converge: $\Xi_{c,K}(s)\to\Xi_{c}(s)$ for each $s$, by dominated convergence of the product factors. Taking $K\to\infty$ in (7):
Step 3 (Contour integration). For a single component $E_{k}(u)=-\frac{1}{\pi}\alpha_{k}^{-1/2}\sin(\omega_{k}u+\phi_{k})$, we evaluate $I_{k}=\int_{0}^{\infty}\frac{2w^{2}\sin(\omega_{k}u+\phi_{k})}{u(u^{2}+w^{2})}\,du$ by contour integration. (The lower limit in (7) is $\gamma_{1}^{(0)}\approx 17.8$, not 0; the difference $\int_{0}^{\gamma_{1}^{(0)}}$ is bounded and absorbed into the $O(c)$ term.)
Write $\frac{2w^{2}}{u(u^{2}+w^{2})}=\frac{2}{u}-\frac{2u}{u^{2}+w^{2}}$ and use the standard integrals (for $\omega>0$, $\operatorname{Re}(w)>0$):
$\int_{0}^{\infty}\frac{\sin(\omega u+\phi)}{u}\,du$ $=\frac{\pi}{2}\cos\phi+\text{(Ci terms)},$
$\int_{0}^{\infty}\frac{u\sin(\omega u+\phi)}{u^{2}+w^{2}}\,du$ $=\frac{\pi}{2}\operatorname{Re}\left[e^{i\phi}\,e^{-\omega w}\right]+\text{(Ci terms)}.$
The Ci (cosine integral) terms cancel in the difference. The key identity, verified by residue calculus at the poles $u=\pm iw$ of $1/(u^{2}+w^{2})$, is:
@@
(8) $I_{k}=\pi\cos\phi_{k}\cdot\bigl(1-e^{-\omega_{k}\operatorname{Re}(w)}\cos(\omega_{k}\operatorname{Im}(w))\bigr)+\pi\sin\phi_{k}\cdot e^{-\omega_{k}\operatorname{Re}(w)}\sin(\omega_{k}\operatorname{Im}(w)).$
For $s=\sigma+it$: $\operatorname{Re}(w)=\sigma-1/2$, $\operatorname{Im}(w)=t$. The constant term ($1$ inside the parenthesis) gives a $t$-independent contribution: multiplied by $c/(\pi\alpha_{k}^{1/2})$ and summed, this produces $c\sum_{k}\cos(\phi_{k})/\alpha_{k}^{1/2}$, a convergent sum (by Lemma 2.1 with $T=0$) of magnitude $O(c)$. The oscillating terms combine to:
$I_{k}^{\mathrm{osc}}=\pi\,e^{-\omega_{k}(\sigma-1/2)}\bigl[-\cos\phi_{k}\cos(\omega_{k}t)+\sin\phi_{k}\sin(\omega_{k}t)\bigr]=-\pi\,\alpha_{k}^{-(\sigma-1/2)}\cos(\omega_{k}t+\phi_{k}),$
using $e^{-\omega_{k}(\sigma-1/2)}=\alpha_{k}^{-(\sigma-1/2)}$ and the angle-addition identity. Substituting into (7) with $E_{K}=-\frac{1}{\pi}\sum_{k=1}^{K}\alpha_{k}^{-1/2}\sin(\omega_{k}u+\phi_{k})$, the contribution of the $k$-th term is $c\,\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})$ (the factors of $\pi$ cancel). Summing over $k=1,\ldots,K$:
$\operatorname{Re}\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\sum_{k=1}^{K}\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c).$
The $O(c)$ error has two sources: the constant Fourier terms $c\sum_{k}\cos(\phi_{k})/\alpha_{k}^{1/2}$ (a convergent sum independent of $s$) and the lower-limit correction $c\int_{0}^{\gamma_{1}^{(0)}}(\cdots)\,du$ (bounded uniformly for $\sigma\geq 1$, $|t|\geq 2$). Both are independent of $K$; the total error is $c\cdot C_{0}+O(c^{2})$ where $C_{0}=C_{0}(\theta)$ is an explicit constant depending on the phase parameter $\theta$.
Taking $K\to\infty$: the LHS converges to $\operatorname{Re}\log(\Xi_{c}(s)/\Xi_{0}(s))$ (since $\Xi_{c,K}\to\Xi_{c}$ pointwise), and the trigonometric sum on the RHS converges for $\sigma\geq 1$ (absolutely for $\sigma>1$; conditionally at $\sigma=1$ by Lemma 2.1). The linearization error $\sum\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2}$ for the full series converges by Lemma 1.2, contributing $O(c^{2})$. The identity (6) follows with total error $O(c)$. ∎

### v2

@@ Abstract.
We construct a single explicit entire function $\Xi_{c}(s)$ of order 1, with all zeros provably on $\operatorname{Re}(s)=1/2$, satisfying a functional equation $\Xi_{c}(s)=\Xi_{c}(1-s)$, whose normalized form $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$ is uniformly bounded for $\operatorname{Re}(s)>1+\delta$ (any $\delta>0$) yet satisfies $\sup_{t}|Z_{c}(1+it)|=+\infty$. The function thus satisfies an analogue of the Riemann Hypothesis together with the sharp bounded/unbounded transition at $\sigma=1$ characteristic of $\zeta$. The transition is controlled by a Dirichlet series $D(s)=\sum e^{-ik\theta}\,p_{k}^{-s}$ whose absolute convergence for $\sigma>1$ and divergence at $\sigma=1$ drive the dichotomy, in direct analogy with $\log\zeta(s)=\sum p^{-s}+O(1)$. The key technical input is a dyadic large-sieve estimate establishing the linearization condition that connects the Hadamard product to $D$. The construction and proofs were developed in collaboration with Claude (Anthropic); see Acknowledgments.
@@ 1. Introduction
  • On $\sigma=1$: $\limsup_{t\to\infty}|\zeta(1+it)|=\infty$, with $|\zeta(1+it)|=\Omega(\log\log t)$ [17, Thm. 8.5].
@@
The mechanism behind $\zeta$’s transition is well understood at the level of Dirichlet series: $\log\zeta(s)=P(s)+O(1)$ for $\sigma>1$, where $P(s)=\sum p^{-s}$ is the prime zeta function [17, §1.6, §3.2]. Boundedness for $\sigma>1$ follows from the absolute convergence of $\sum p^{-\sigma}$; unboundedness on $\sigma=1$ from its divergence (Mertens [14, Thm. 2.7]). The zero-counting error $S(T)=\pi^{-1}\arg\zeta(\tfrac{1}{2}+iT)$ satisfies $S(T)\approx-\pi^{-1}\sum_{p}\sin(T\log p)/\sqrt{p}$ in $L^{2}$ [17, Thm. 14.22]—the imaginary part of $P(\tfrac{1}{2}+iT)/\pi$.
Our construction replaces $P(s)$ with a twisted variant $D(s)=\sum e^{-ik\theta}\,p_{k}^{-s}$ (with $\theta\in(0,2\pi)$), uses its restriction to $\sigma=1/2$ to perturb the zeros of a Hadamard product, and proves that the resulting function’s log-ratio with the unperturbed product is approximately cD(s)-cD(s)—the “spectral identity” of this paper.
@@
Fix $\theta\in(0,2\pi)$ (e.g., $\theta=1$) and $c>0$ sufficiently small. Define $\Xi_{c}(s)$ as in §2, with $Z_{c}(s)=\Xi_{c}(s)/[\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)]$. Then:
@@
    For each $\delta>0$, $\sup_{t}|Z_{c}(1+\delta+it)|\leq\exp\bigl(c\,P(1+\delta)+O(1)\bigr)$, where $P(\sigma)=\sum_{p}p^{-\sigma}$. Moreover $P(\sigma)\to 0$ as $\sigma\to\infty$, so the bound tends to $e^{O(1)}$: the ratio $\Xi_{c}/\Xi_{0}$ is uniformly close to a constant for $\sigma$ large.
@@
  6. (f)
    The ratio $\Xi_{c}/\Xi_{0}$ admits the approximate Euler product $\Xi_{c}(s)/\Xi_{0}(s)=\kappa_{0}\cdot\prod_{k}\exp(-c\,e^{-ik\theta}p_{k}^{-s})\cdot M(s)$ for $\sigma>1$, where $\kappa_{0}\neq 0$ is an explicit constant, $M(s)$ is bounded (with $M\to 1$ as $c\to 0$), and the Euler product is itself a Dirichlet series with multiplicative coefficients.
2. The Dirichlet Series and Construction
2.1. The Dirichlet series $D(s)$
@@
Fix $\theta\in(0,2\pi)$ (e.g., $\theta=1$), an amplitude $c>0$, and let $p_{1}<p_{2}<\cdots$ denote the primes. Define the Dirichlet series
(1) $D(s)=\sum_{k=1}^{\infty}e^{-ik\theta}\,p_{k}^{-s},$
which converges absolutely for $\sigma>1$ (since $\sum p_{k}^{-\sigma}<\infty$) and conditionally for $\sigma>0$, by Abel summation: the partial sums $G_{K}=\sum_{k=1}^{K}e^{-ik\theta}$ satisfy $|G_{K}|\leq 2/|1-e^{-i\theta}|=:C_{\theta}$ (finite since $\theta\in(0,2\pi)$), and the amplitudes $p_{k}^{-\sigma}$ decrease monotonically to zero. Viewing $D$ as an ordinary Dirichlet series $\sum b_{n}\,n^{-s}$ (with $b_{n}=e^{-ik\theta}$ for $n=p_{k}$, zero otherwise), the partial sums $B(x)=\sum_{n\leq x}b_{n}$ satisfy $|B(x)|\leq C_{\theta}$, so the abscissa of convergence is $\sigma_{c}=0$ (cf. [14, Thm. 1.3]). In the strip of conditional convergence, [14, Thm. 1.5] gives
(2) $|D(\sigma+it)|\ll(|t|+4)^{1-\sigma+\varepsilon}$
for any fixed $\sigma\in(0,1)$ and $\varepsilon>0$. For $\sigma>1$ the stronger bound $|D(\sigma+it)|\leq P(\sigma)=\sum p^{-\sigma}$ holds by absolute convergence.
Define the perturbation function
(3) $E(T)=\frac{1}{\pi}\,\operatorname{Im}\bigl[D(\tfrac{1}{2}+iT)\bigr]=-\frac{1}{\pi}\sum_{k=1}^{\infty}\frac{\sin(T\log p_{k}+\phi_{k})}{p_{k}^{1/2}},$
where $\phi_{k}=k\theta$. The convergence of $E(T)$ for each $T>0$ follows from that of $D(1/2+iT)$, and by (2), $|E(T)|=O(|T|^{1/2+\varepsilon})$. We write $\omega_{k}=\log p_{k}$ for brevity in Fourier-analytic arguments.
@@ 2.2. The perturbed zeros
and let $\Xi_{c}(s)$ be the Hadamard product (cf. [6, §1.10])
(6) $\Xi_{c}(s)=\prod_{n=1}^{\infty}\left(1+\frac{(s-\tfrac{1}{2})^{2}}{\gamma_{n}^{2}}\right).$
Lemma 2.1 (Well-definedness).
For $c$ sufficiently small, $\gamma_{n}>0$ for all $n$, $\sum\gamma_{n}^{-2}<\infty$, and $\Xi_{c}(s)$ is an entire function of order $1$.
@@
Since $|E(T)|=O(|T|^{1/2+\varepsilon})$, $|\delta_{n}|=O(c(\gamma_{n}^{(0)})^{1/2+\varepsilon}/\log\gamma_{n}^{(0)})=o(\gamma_{n}^{(0)})$, so $\gamma_{n}>\gamma_{n}^{(0)}/2$ for all $n$ sufficiently large. For the finitely many small $n$, $\delta_{n}$ is finite (since $E(\gamma_{n}^{(0)})$ converges for each $n$) and proportional to $c$, so taking $c$ small enough ensures $\gamma_{n}>0$ for all $n$. Since $\gamma_{n}^{(0)}\sim 2\pi n/\log n$, $\sum\gamma_{n}^{-2}<\infty$ and the exponent of convergence equals $1$. ∎
We set $Z_{c}(s)=\Xi_{c}(s)/\Gamma_{\pi}(s)$ and $Z_{0}(s)=\Xi_{0}(s)/\Gamma_{\pi}(s)$, where $\Gamma_{\pi}(s)=\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)$. Note that $Z_{c}/Z_{0}=\Xi_{c}/\Xi_{0}$ (the $\Gamma_{\pi}$ factors cancel).
@@ 3. The Spectral Identity
The central result relates the Hadamard product ratio $\Xi_{c}/\Xi_{0}$ to the Dirichlet series $D$.
@@
Let $\omega_{k}=\log p_{k}$ and $\phi_{k}=k\theta$ as in §2. For $\sigma>1/2$ and $|t|\geq 2$:
(7) $\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}}(\sigma+it)=-c\sum_{k}p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(1).$
Equivalently, there exists a bounded analytic function $R(s)$ with $R(s)=O(1)$ uniformly for $\sigma>1/2$, $|t|\geq 2$, such that
(8) $\log\frac{\Xi_{c}}{\Xi_{0}}(s)=-c\,D(s)+R(s).$
The cosine sum converges absolutely for $\sigma>1$ and conditionally for $\sigma>0$ (as $\operatorname{Re}\,D(\sigma+it)$).
@@
We establish (7) first, then deduce (8) by analyticity. Throughout, $c$ is fixed, so all implicit constants may depend on $c$ and $\theta$.
Step 1 (Linearization). By (5), $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$. Each Hadamard factor satisfies:
@@
Step 2 (Sum to integral). For each $K$, let $\Xi_{c,K}$ be the Hadamard product with zeros perturbed using $E_{K}$ (first $K$ terms of $E$). Since $|E_{K}|\leq\frac{1}{\pi}\sum_{k=1}^{K}p_{k}^{-1/2}$ (a constant for each $K$), the perturbations are $\delta_{n,K}=O(1/\log\gamma_{n}^{(0)})$, giving convergent quadratic remainder $\sum\delta_{n,K}^{2}/(\gamma_{n}^{(0)})^{2}<\infty$. The linearized sum is a Riemann sum for $\int\frac{2w^{2}\,cE_{K}(u)}{u(u^{2}+w^{2})}\,du$ (one term per zero, spacing $\sim 1/N_{\mathrm{main}}^{\prime}$). The integrand is smooth for $u>0$ and $\sigma>1/2$ (the poles of $1/(u^{2}+w^{2})$ lie off the real axis), so by Euler–Maclaurin:
(9) $\operatorname{Re}\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\int_{\gamma_{1}^{(0)}}^{\infty}\operatorname{Re}\frac{2w^{2}\,E_{K}(u)}{u(u^{2}+w^{2})}\,du+O(1).$
Step 3 (Evaluation of the integral). Using the partial fraction $\frac{2w^{2}}{u(u^{2}+w^{2})}=\frac{2}{u}-\frac{2u}{u^{2}+w^{2}}$ and the standard integrals $\int_{0}^{\infty}\frac{\sin\omega u}{u}\,du=\frac{\pi}{2}$ (Dirichlet) and $\int_{0}^{\infty}\frac{u\sin\omega u}{u^{2}+\beta^{2}}\,du=\frac{\pi}{2}e^{-\omega\beta}$ for $\omega>0$, $\operatorname{Re}\beta>0$ ([8, §3.723], extended to complex $\beta$ by analytic continuation), the contribution of each component $\sin(\omega_{k}u+\phi_{k})$ decomposes into convergent pieces. [^1] The individual integrals $\int\sin(\omega u+\phi)/u\,du$ and $\int u\sin(\omega u+\phi)/(u^{2}+w^{2})\,du$ each diverge at $u=0$ when $\sin\phi\neq 0$ (from the $\cos(\omega u)/u$ term), but the divergent parts cancel in the partial-fraction difference. The oscillatory contribution of the $k$-th component is $\pi\,p_{k}^{-(\sigma-1/2)}\cos(\omega_{k}t+\phi_{k})$ (using $e^{-\omega_{k}(\sigma-1/2)}=p_{k}^{-(\sigma-1/2)}$ and the angle-addition identity); the non-oscillatory part $\pi\cos\phi_{k}(1-p_{k}^{-(\sigma-1/2)})$ sums to a bounded quantity (convergent by Abel summation, depending on $\sigma$ but not on $t$).
Substituting into (9) with the amplitudes $1/(\pi p_{k}^{1/2})$ from $E_{K}$, the $k$-th term contributes $c\,p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})$. Including the bounded lower-limit correction $\int_{0}^{\gamma_{1}^{(0)}}(\cdots)\,du$:
@@
(10) $\operatorname{Re}\log\frac{\Xi_{c,K}(s)}{\Xi_{0}(s)}=-c\sum_{k=1}^{K}p_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(1)+O\Bigl(\sum_{n}\delta_{n,K}^{2}/(\gamma_{n}^{(0)})^{2}\Bigr),$
where the $O(1)$ accounts for the non-oscillatory Fourier terms, the lower-limit correction, and the Euler–Maclaurin error (all independent of $K$), while the final term is the quadratic Taylor remainder from Step 1. For each fixed $K$ this remainder is finite, but it grows with $K$ (since $M_{K}$ grows).
Taking $K\to\infty$: the cosine sum converges (absolutely for $\sigma>1$; conditionally for $\sigma>0$), and $\Xi_{c,K}\to\Xi_{c}$ (since the zero sequences converge). The only obstruction to passing (10) to the limit is the Taylor remainder: for the full (non-truncated) series, $\delta_{n}=-cE(\gamma_{n}^{(0)})/N_{\mathrm{main}}^{\prime}(\gamma_{n}^{(0)})$ and $|E(\gamma_{n}^{(0)})|$ is no longer bounded by a fixed constant (its RMS grows like $\sqrt{\log\log n}$ by (14)). The quadratic remainder $\sum\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2}$ converges iff $\sum E(\gamma_{n}^{(0)})^{2}/n^{2}<\infty$, which is the content of Lemma 3.2 below. This establishes (7).
Step 4 (Complex extension). Equation (7) controls the real part of $\log(\Xi_{c}/\Xi_{0})+cD$; we now upgrade to the full complex identity. Define $F(s)=\log(\Xi_{c}(s)/\Xi_{0}(s))+cD(s)$. Both terms are analytic for $\sigma>1/2$: the Hadamard ratio $\Xi_{c}/\Xi_{0}$ is nonvanishing there (all zeros of $\Xi_{0}$ and $\Xi_{c}$ lie on $\sigma=1/2$), and $D$ converges for $\sigma>0$. By (7) (proved above without reference to (8)), $\operatorname{Re}F(s)=O(1)$ for $\sigma>1/2$, $|t|\geq 2$. Evaluating at $s=2$: $D(2)$ converges absolutely, and $\log(\Xi_{c}(2)/\Xi_{0}(2))$ is finite (each Hadamard factor at $s=2$ differs from $1$ by $O(\delta_{n}/\gamma_{n}^{2})$, and $\sum|\delta_{n}|/\gamma_{n}^{2}<\infty$), so $F(2)=O(1)$. Since $F$ is analytic with bounded real part, the Borel–Carathéodory theorem gives $|F(s)|=O(1)$ for $\sigma>1/2$. Setting $R(s)=F(s)$ gives (8). ∎
3.1. Linearization convergence
The following lemma ensures that the Taylor linearization in Step 1 has convergent remainder for the full (non-truncated) series, which is the key input for passing the spectral identity from finite truncations to $\Xi_{c}$ itself.